OpenStudy (sara4954):

A stone is thrown vertically upward with a speed of 27.0 m/s .How fast is it moving when it is at a height of 14.0 m? How much time is required to reach this height?

OpenStudy (mathmate):

Since this is a post in physics, we'll use the usual kinematics equation. v^2-u^2=2aS v=final velocity u=initial velocity a=acceleration S=distance Substitute numbers, v=14 m/s (going up) u=27 m/s (going up) a=-9.81 m/s^2 (going down) S=(v^2-u^2)/2a=\(14^2-27^2)/(-9.81)=? (The answer should be positive) Note that v could be negative, i.e. v=-14 m (going down). But the answers are the same, since v is squared. For the second part, use v(t)=v(0)+at where a=-9.81 v(t)=14 m/s v(0)=27 m/s solve for t. "How much time is required to reach this height" implies time required while going UP, so v(t)=-14 m/s (going down) can be discarded. As you can see, once the basic equations of kinematics have been mastered (or even just memorized), solutions follow.

OpenStudy (sachintha):

mathmate you have misinterpreted the question. For the first part it asks for the velocity at the height of 14m. So, \(v^2-u^2=2as\) \(\,\qquad v_a = \sqrt{2as+u^2}\) \(\:\,\quad\qquad = \sqrt{2\times-9.8\times14+27^2}\) For the second part, \(v=u+at\) \(\,t = \Large\frac{v-u}{a}\) \(\:\:\;\, = \Large\frac{v_a-27}{-9.8}\) Simplify to get the answers.

OpenStudy (mathmate):

Very true. I missed that. Thank you for point it out. @sachintha